OpenStudy (anonymous):

Been having troubles with this one Juan and Maria tranplanted their sunflowers. Juan's plant is 7.2 cm tall to begin with and is expected to grow 0.6 cm per day. Maria's plant is 4.8 cm tall to begin with and will grow 0.8 cm per day. A. Write two equations that represent the relationship between height (y) and numbers of days (x) for Juan and Maria's sunflowers. Show all your work. B. After how many days will these plants be the same height? Show all your work. C. How tall will the plants be when they are the same height? Show all your work.

OpenStudy (anonymous):

A y= 0.6x+7.2 (for Juan)-------(1 y= 0.8x+4.8 (for Maria) -----(2 B) from 1 and 2 we have 0.6x+7.2 = 0.8x+4.8 i.e. 7.2-4.8= 0.8x-0.6x i.e. 2.4 =0.2x x= 2.4/0.2= 24/2= 12 Hence after 12 days the height of both the plants will be same C) substituting x= 12 in any of the equations say in eq 1) we find y= 0.6x+7.2= 0.6*12+7.2=7.2+7.2 = 14.4 cm Hence after attaining the same height plants will be of 14.4 cm high

OpenStudy (anonymous):

thank you very much you got a new fan